# Compositional Difference-in-Differences for categorical outcomes

### v2

@@ Abstract.
In difference-in-differences (DiD) settings with categorical outcomes, such as voting, occupation, or major choices, treatments often affect both total counts (e.g., turnout) and category shares (e.g., vote shares). Traditional linear DiD models can yield invalid counterfactuals in this context (e.g. negative values ) and lack compatibility with standard discrete choice models. I propose Compositional DiD (CoDiD), which identifies counterfactual totals and shares under a parallel growths assumption: absent treatment, each category’s size grows or shrinks at the same proportional rate in treated and control groups. I show that under a random utility model, this is equivalent to parallel trends in expected utilities, i.e., the change in average latent desirability for each alternative is identical across groups. Consequently, relative preferences (e.g., how individuals prefer Democrat vs. Republican) evolve in parallel, and counterfactual distributions follow parallel trajectories in the probability simplex. I extend CoDiD to (i) derive bounds under relaxed assumptions, (ii) accommodate staggered treatment timing, and (iii) construct a synthetic DiD analog. I illustrate the method’s empirical relevance through two applications: first, I examine how early voting reforms affect voter choice in U.S. presidential elections; second, I analyze how the Regional Greenhouse Gas Initiative (RGGI) affected the electricity mix in participating states.
Keywords: categorical outcomes, difference-in-differences, causal inference, compositional data, voting, rggi.
@@
Corresponding address: oib5044@psu.edu, Department of Economics, The Pennsylvania State University, University Park, PA 16802. I am grateful to Marc Henry, Andres Aradillas-Lopez, Sung Jae Jun, and Keisuke Hirano for their guidance on this project. I am also grateful to the participants of the Penn State Econometrics Seminar, the Penn State Student Econometrics Workshop, and the RESA Alumni Conference 2025 for their valuable comments and suggestions. All errors are mine.
@@ Introduction
The classical difference-in-differences (DiD) framework identifies the average treatment effect on the treated under the parallel trends assumption for scalar outcomes. However, in many applications, the outcome of interest is categorical, and the notion of an average level is generally undefined. Examples include voting decision (democrat, republican, others), occupational sector (agriculture, manufacturing, and services) or field of study (stem, humanities, business, or others). For categorical outcomes, the key quantitative objects are the category-specific quantities, which jointly determine both the overall total quantity and the corresponding probability mass function across categories. In this case, one would like to analyze both absolute changes (how many units are added or removed in absolute terms for each category, and overall) and relative changes (how the shares are reallocated across categories) in a consistent way. For example, a voting policy can differently affect both the total number of voters (turnout) and the reallocation of vote shares for each party. A minimum wage policy can change both the number of people across all employment sectors and the probability of being in each of them. An education reform may reallocate students between fields, but policymakers may also care about the total number of new graduates overall.
The classical linear parallel trends assumption, which imposes a linear evolution in outcome levels, may be inappropriate in this context for several reasons already noted in the literature (see Puhani (2012); Wooldridge (2023)). A naive strategy would be to apply linear parallel trends separately to the raw quantities of each category, obtain counterfactual quantities, and then normalize them to form a counterfactual distribution. However, this approach is problematic for three main reasons. First, applying linear parallel trends to raw quantities is scale-dependent, as it ignores differences in group size. Transferring the same additive change from a large control group to a smaller treated group can distort counterfactual quantities, exaggerating or understating effects, and thus producing misleading totals and distributions. Second, this method can yield negative counterfactual quantities when a negative shift in the control group outweighs the pre-treatment value in the treated group. Finally, it would imply choice dynamics that lack any behavioral or structural justification from economic theory regarding how choices across categories evolve.
On top of that, applying parallel trends directly to probabilities is theoretically possible but often yields invalid counterfactuals, with values below zero or exceeding one. The resulting counterfactual distribution can fall outside the probability simplex (the set of all vectors of positive numbers summing to one), rendering them invalid (see figure 1). Most existing nonlinear methods for categorical distributions require a natural ordering of the support because they rely on cumulative distribution functions, which are undefined for unordered outcomes. For such outcomes, only the probability mass function is well-defined, and importantly, these methods typically do not directly provide counterfactuals for the total quantities.
(1,0,0)(1,0,0)(0,1,0)(0,1,0)(0,0,1)(0,0,1)Treated ($t=0$)counterfactual ($t=1$)Control ($t=0$)Control ($t=1$)
Figure 1. Illustration of how parallel trends can lead to invalid counterfactuals in the 2-dimensional probability simplex. The control group moves from $(0.7,0.2,0.1)$ to $(0.3,0.3,0.4)$, while applying the same linear shift to the treated group at $(0.2,0.3,0.5)$ produces an invalid counterfactual (0.2,0.4,0.8)(-0.2,0.4,0.8), shown above as lying outside the probability space.
Ideally, one would like to consider a nonlinear parallel trends assumption on categorical quantities that has a clear natural meaning, resolves the scale-dependence problem, and ensures that the evolution of shares remains consistent with standard discrete choice theory. This paper introduces Compositional difference-in-differences (CoDiD), a new framework for analyzing categorical outcomes in a DiD setting. The method is built on a nonlinear parallel trends assumption, which I call parallel growths, which allows us to point-identify the counterfactual categorical quantity, and thus the counterfactual total and probability mass functions within a DiD framework. I do so while addressing all the challenges outlined above. In addition, I define two treatment effect measures tailored to this context. The growth treatment effect on the treated (GTT) quantifies how each category’s absolute size changes as a result of treatment, and also how the total quantity has evolved as a result of treatment. The second is the compositional treatment effect on the treated (CTT), which captures how the probability mass is reallocated across categories in response to treatment.
I begin with the canonical $2\times 2$ DiD setting. The parallel growths assumption is a parallel trends of log-transformed categorical quantities, meaning that the absolute size of each category would grow or shrink at the same proportional rate in both treated and control groups. Using a standard random utility model, I show that this assumption is equivalent to assuming parallel trends of latent utilities for each category. This means that in the absence of treatment, both groups would have experienced the same evolution in how desirable each option is on average. For example, if in the control group, the expected utility for voting democrat increase, the same movement will happen in treated group in absence of treatment. The implied evolution of the shares is a parallel trends in the log-odds of category shares. In the random utility model, log-odds capture utiliy differences, therefore relative preferences. Under this interpretation, parallel growths mean that, absent treatment, the evolution of relative preferences between any pair of categories would be similar in both groups. In other words, the evolution of the way people chose between categories in the control group will be similar to what happens in the treated group. Under this interpretation, we can view the causal effect as the way the treatment shift the way categories are desired.
Furthermore, without relying on a random utility framework, I derive another economic implication of parallel growths. In the control group, category shares naturally evolve over time, with some categories gaining mass at the expense of others. The assumption implies that, in the absence of treatment, the treatment group would have experienced the same reallocation pattern observed in the control group. This interpretation allows us to view the causal effect as the extent to which the treatment change the way probability mass is redistributed across categories. Finally, in geometric terms, each probability mass function (PMF) corresponds to a point in the simplex, and its evolution traces a path within this space. The parallel growths assumption means that, in the absence of treatment, the treated group’s counterfactual path would be parallel to the control group’s path under Aitchison geometry, the standard geometry for the simplex in compositional data analysis (Aitchison (1982); Egozcue et al. (2003)).
I also discuss the connections and key differences between my approach and the multinomial logistic two-way fixed effects model. In the simple $2\times 2$ case, I show that, when focusing solely on category shares, both methods yield the same counterfactual shares. The main distinction lies in the fact that my approach also recovers counterfactual quantities, which cannot be obtained directly from the logistic model. Moreover, my framework introduces a new treatment effect parameter that offers a more natural and interpretable measure than the change in log-odds implied by the logistic specification. Even when the analysis is limited to shares, the differences become more pronounced when extending the framework to multiple time periods (e.g. staggered adoption case) and relaxations of the identifying assumption. My approach provides a more flexible and easily extendable foundation for such generalizations. Another comparison is made with the DiD and the changes-in-changes (CiC) frameworks. While DiD targets on expectations (points on the real line) and CiC, cumulative distribution functions (points in a space of CDFs), my method also targets probability mass functions (points in the probability simplex). The underlying identification logic is consistent across all three methods: each characterizes the control group’s evolution via a translation mapping and applies that same mapping to the treated group’s initial state to recover its counterfactual. The fundamental distinction, therefore, lies not in the core logic but in the object being transported, expectations, CDFs, or PMFs, and the geometry that defines the translation. From this unified perspective, the CoDiD emerges naturally as the DiD/CiC analogue for probability mass functions.
In a more general framework with multiple time periods, the parallel growths assumption can be assessed visually by plotting the log-transformed quantities. If the trends appear non-parallel in pre-treatment periods, the assumption becomes less credible. In this case, I first propose a set of relaxations of the assumption in the spirit of Ban and Kédagni (2022). Specifically, I bound the counterfactual log differences between groups by restricting their change to lie within the convex hull of observed changes across pre-treatment periods. This approach produces bounds for both counterfactual quantities and counterfactual distributions. When multiple potential control groups are available, I develop a synthetic DiD (see Arkhangelsky et al. (2021)) analog, synthetic CoDiD, that merges the flexibility of synthetic control methods with the causal logic of difference-in-differences. This extension constructs a weighted combination of control units whose pre-treatment log-trajectory closely matches that of the treated unit, while preserving the parallel growths structure in the post-treatment period. The resulting control group not only improves pre-treatment fit but also maintains interpretability within the CoDiD framework. Finally, I generalize the methodology to settings with staggered treatment adoption, where units receive treatment at different points in time. I follow the same assumptions as in Callaway and Sant’Anna (2021), propose two types of parallel growths assumption in this case, and derive the corresponding counterfactual in each case.
As applications, I first examine the impact of early voting programs on the distribution of voter support across political parties in U.S. presidential elections. These programs, allowing voters to cast ballots before election day, either in person or by mail, have been the subject of intense partisan debate. I revisit this question by analyzing both total turnout and the compositional shift in voter preferences. I consider Maryland and New Jersey, which introduced early voting before 2008 in treated group, while Pennsylvania and New York serve as the control group. As expected, the treatment increased overall turnout ( increase of $4.4\%$) and the number of votes for all parties. However, the compositional treatment effect on the Treated (CTT) reveals a nuanced reallocation: a previously indifferent voter (assigning equal initial weight to all parties) would, after the treatment, shift support more toward alternative parties, with both major parties losing relative appeal, Republicans more so than Democrats. Secondly, I analyze the Regional Greenhouse Gas Initiative (RGGI) and its effects on the mix of electricity generation sources in participating states. RGGI, a cap-and-trade program designed to reduce carbon emissions from the power sector, is expected to shift the composition of electricity generation across coal, natural gas, nuclear, and renewable sources by promoting cleaner technologies and discouraging carbon-intensive ones. I assess how RGGI affected the size of each source in the electricity generation process over time. I find that RGGI sharply cut fossil fuel use, especially coal, by shifting generation to gas and nuclear rather than boosting renewables, underscoring the need for complementary policies to accelerate a truly renewable-driven energy transition.
@@
First, it contributes to the extensive DiD literature. Classical DiD focuses on scalar outcomes, estimating treatment effects on the mean, with surveys and applications summarized in Lechner (2011); Cafri et al. (2019); De Chaisemartin and d’Haultfoeuille (2023); Roth et al. (2023); Baker et al. (2025). Recent works have extended DiD to more complex settings: Callaway and Sant’Anna (2020) develops methods for staggered treatment adoption, Arkhangelsky et al. (2021) combines DiD with synthetic controls, and Manski and Pepper (2018); Rambachan and Roth (2020); Ban and Kédagni (2022) relax the parallel trends assumption. A second wave of research extends DiD to entire outcome distributions. Early contributions include quantile DiD Meyer et al. (1990) and the changes-in-changes (CiC) approach Athey and Imbens (2006), which identifies counterfactual distributions non-parametrically. Subsequent work has generalized these ideas to multivariate outcomes Torous et al. (2021), settings where identifying assumptions apply to cumulative distribution functions Havnes and Mogstad (2015); Roth and Sant’Anna (2021), copula-based approaches Callaway et al. (2018); Callaway and Li (2019); Ghanem et al. (2023), and methods using characteristic functions Bonhomme and Sauder (2011). However, these methods do not directly address categorical outcome distributions, which are probability mass functions. Graves et al. (2022) proposes a DiD for categorical outcomes while relying on parallel trends for proportions, which I explained earlier is not appropriate in general. Zhou et al. (2025) extend the difference-in-differences framework to non-Euclidean data, including probability mass functions, by employing Fréchet means and geodesic transport. While their approach is mathematically elegant, the identifying assumptions are primarily geometric and offer limited economic interpretability. In contrast, my framework delivers both counterfactual quantities and probability mass functions that retain economic intuition while incorporating a geometric perspective specifically designed for categorical outcomes. Another related paper is Tchetgen Tchetgen et al. (2024), who propose a general DiD framework applicable to count data in the binomial setting. My analysis differs by focusing on the multinomial case, allowing for richer categorical structures and more general forms of distributional change. Second, the paper contributes to the compositional data analysis (CoDA) literature. CoDA, pioneered by Aitchison (1982, 1990, 1992, 2002), provides tools for analyzing data constrained to the simplex. Subsequent developments Egozcue et al. (2003); Billheimer et al. (2001); Barceló-Vidal et al. (2001) refine transformations and models that respect the geometric structure of compositional data. More recently, Arnold et al. (2020) connected CoDA to causal inference. I also combine compositional data methods with econometric identification strategies to analyze causal effects when outcomes are probability mass functions, providing a bridge between DiD and compositional statistics.
@@
The paper is organized as follows. Section 1 introduces the canonical $2\times 2$ case, presenting the main identifying assumption along with its economic and geometric interpretations. Section 2 generalizes the framework to settings with multiple time periods. Section 3 discusses the connection with existing approaches, and Section 4 illustrates the methodology through two empirical applications. Finally, Section 5 concludes.
@@ 1. Canonical $2\times 2$ DiD Case
In this section, I focus on the canonical, two-group, two-time period case in the DiD framework, and I clearly expose the assumptions and their implications. I will start with the analytical framework, later discuss the main identifying assumption, and all the economic and geometric implications.
1.1. Parallel growths and Identification
1.1.1. Analytical framework
In this section, I introduce the basic framework of my method. A categorical outcome $Y$ takes values in a finite set of $p$ mutually exclusive categories:
$Y\in\{c_{1},c_{2},\dots,c_{p}\}.$
I focus primarily on the absolute quantities for each category, which can appear in several empirical contexts. They can arise when the outcome is observed at the individual level. For example, in evaluating a job training program, individuals may be classified as unemployed, part-time employed, or full-time employed. The categorical quantity is obtained by aggregating individuals into these outcome categories. Sometimes, the data consist directly of quantities across categories rather than individual-level observations. For example, the number of votes each political party receives in an election. Unlike individual-level data, the identities or attributes of the units are not observed. In all these cases, the proposed method remains applicable, as it relies solely on knowledge of the categorical quantities.
@@ 1.1.1. Analytical framework
I consider a setting with two groups, $g\in\{0,1\}$, where $g=1$ is the treated group and $g=0$ is the control group. These groups are observed in two time periods, $t\in\{0,1\}$, with $t=0$ being the pre-treatment period and $t=1$ the post-treatment period. The object of interest is a categorical outcome. For any category $c_{k}$, I define $q^{N}_{g,t}(c_{k})$: the potential untreated quantity (the quantity in $c_{k}$ if the group $g$ were untreated at time $t$). I also have $q^{I}_{g,t}(c_{k})$: the potential treated quantity (the quantity in $c_{k}$ if the group $g$ were treated at time $t$). For the untreated potential outcomes, I aggregate across categories to define the total quantity for a group-period:
$S^{N}_{gt}=\sum_{k=1}^{p}q^{N}_{g,t}(c_{k}).$
The share (or probability) for category $c_{k}$ is then:
$\pi^{N}_{g,t}(c_{k})=\frac{q^{N}_{g,t}(c_{k})}{S^{N}_{gt}}.$
The vector of quantities and the entire probability mass function are represented as:
$q^{N}_{g,t}=\left[q^{N}_{g,t}(c_{1}),\dots,q^{N}_{g,t}(c_{p})\right],\quad\pi^{N}_{g,t}=\left[\pi^{N}_{g,t}(c_{1}),\dots,\pi^{N}_{g,t}(c_{p})\right],\quad\text{where}\quad\sum_{k=1}^{p}\pi^{N}_{g,t}(c_{k})=1.$
@@
All the same definitions apply to the treated potential quantities, $q^{I}_{g,t}(c_{k})$, and their associated totals and shares. From the data, I can identify the vectors $q^{N}_{0,0}$ for the control group and $q^{N}_{1,0}$ for the treated group in the pre-treatment period. In the post-treatment period, I can identify from data the untreated vector of quantities for the control group, $q^{N}_{0,1}$, and the treated counterpart for the treated group, $q^{I}_{1,1}$. The counterfactual of interest is the vector of quantities that I would have observed for the treated group in the absence of treatment $q^{N}_{1,1}$. Identifying this missing object is the core challenge of my analysis. Once this is identified, I can consider the treatment effect parameters.
@@
I introduce two treatment effect parameters to capture the impact of the intervention. The first is the growth treatment effect on the treated (GTT), which quantifies the causal effect on the absolute size of a category. For each category $c_{k}$ (where $k=1,\ldots,q$), the GTT is defined as the proportional change in its quantity:
$\text{GTT}(c_{k})=\frac{q^{I}_{1,1}(c_{k})}{q^{N}_{1,1}(c_{k})}-1,\quad GTT=\frac{S^{I}_{1,1}}{S^{N}_{1,1}}-1.$
@@
This parameter has an intuitive interpretation: $GTT=0$ implies the treatment had no effect on the absolute. GTT>0GTT>0indicates that the treatment caused category $c_{k}$ to grow in absolute terms. For example, a GTT of $0.15$ means the category’s size increased by $15\%$ due to the treatment. $GTT<0$ signifies that the treatment caused the category to shrink. A GTT of 0.10-0.10, for instance, corresponds to a $10\%$ decline.
@@
The second parameter is the compositional treatment effect on the treated (CTT), which captures how treatment redistributes weight across categories. To measure this, I apply the compositional difference operator commonly used in compositional data analysis to compare probability mass functions. This operator quantifies how each category expands or contracts relative to all others. For intuition, fix a category $c_{k}$ and consider its proportional change:
@@
$r_{k}=\frac{\pi_{1,1}^{I}(c_{k})}{\pi_{1,1}^{N}(c_{k})},\quad\begin{cases}r_{k}>1&\text{category }k\text{ grew},\\ r_{k}<1&\text{category }k\text{ shrunk},\\ r_{k}=1&\text{no change}.\end{cases}$
@@
Dividing by the total change gives a vector showing each category’s relative gain or loss. I obtain the compositional difference operation used in compositional data analysis for the control group.
$\pi_{1,1}^{N}\ominus\pi_{1,1}^{N}=\frac{1}{\sum_{k=1}^{q}r_{k}}\,\big[r_{1},r_{2},\dots,r_{p}\big]=\frac{1}{\sum_{k=1}^{p}\frac{\pi_{1,1}^{I}(c_{k})}{\pi_{0,0}^{N}(c_{k})}}\left[\frac{\pi_{1,1}^{N}(c_{1})}{\pi_{1,1}^{N}(c_{1})},\dots,\frac{\pi_{1,1}^{N}(c_{p})}{\pi_{1,1}^{N}(c_{p})}\right].$ (1)

### v3

@@ Abstract.
Many policy evaluations involve vectors of category-specific quantities, either categorical outcomes (e.g., employment type, major choice) or compositional measures (e.g., GDP by sector, votes by party, electricity generation by source). In these settings, both intensive margins (shares) and extensive margins (totals) can matter. However, existing Difference-in-Differences (DiD) strategies typically focus only on the shares and do not jointly identify treatment effects on totals. In addition, these approaches usually lack a clear economic interpretation. I develop Compositional Difference-in-Differences (CoDiD), a new framework that identifies treatment effects on both shares and totals in a coherent way. The key assumption is parallel growth: in the absence of treatment, the log-quantities of each category would have evolved in parallel for the treated and control groups. I show that, under a random-utility discrete-choice model, this condition is equivalent to parallel trends in expected utilities, meaning that the change in average latent attractiveness for each alternative is identical across groups. Furthermore, geometrically, the counterfactual distributions (shares) follow parallel trajectories in the probability simplex. In settings with multiple time periods, I introduce a relaxation that delivers bounds when parallel growth may not hold. I illustrate the empirical relevance of the method by examining how early voting reforms affected the 2008 U.S. election.
Keywords: categorical outcomes, difference-in-differences, causal inference, compositional data, simplex, voting.
@@
Corresponding address: oib5044@psu.edu, Department of Economics, The Pennsylvania State University, University Park, PA 16801. I am grateful to Marc Henry, Andres Aradillas-Lopez, and Sung Jae Jun for their guidance on this project. I also thank Keisuke Hirano, Ismael Mourifié, Désiré Kédagni, and Patrick Guggenberger, as well as all participants of the Penn State Econometrics Seminar, the Penn State Student Econometrics Workshop, the RESA Alumni Conference 2025, and the EconNect Africa JMC Symposium, for their valuable comments and suggestions. All remaining errors are my own.
@@ Introduction
Many causal questions in Difference-in-Differences (DiD) settings involve vectors of categorical quantities that together form a well-defined total. These arise naturally when individual-level categorical outcomes are aggregated. For instance, in evaluating a minimum wage policy, individuals may be classified as unemployed, part-time employed, or full-time employed, and the compositional vector counts how many fall into each group. In other cases, the data consist directly of quantities across categories, with no underlying individual identifiers, such as the number of votes received by each political party in an election. Compositional vectors also appear when a single quantity is decomposed into meaningful parts, as in electricity generation by source (gas, coal, nuclear, renewable), GDP by sector (agriculture, industry, services), or a budget allocated to each category (for example, food, rent, healthcare, others). In all such cases, the outcome is multidimensional, and both the shares and the total matter for the analysis.
In many empirical settings, researchers focus on studying treatment effect on the intensive margin, captured by the shares of outcomes allocated across categories. When the outcome is categorical, they typically convert unordered qualitative outcomes into binary indicators use parallel trends on each of these binary outcomes. Because shares encode only relative information and are constrained between zero and one, using linear DiD in this case can be problematic, as noted by Lechner (2011); Athey and Imbens (2006); Wooldridge (2023); Ahn and Kasahara (2025)). A range of alternative causal inference methods, such as the geodesic difference-in-differences Zhou et al. (2025), and transition-independence approaches Ahn and Kasahara (2025), are better suited since they cannot generate invalid counterfactuals. These methods allow researchers to trace how the distribution across categories evolves when a policy changes, while respecting the intrinsic constraints of the compositional shares(such as non-negativity and the unit-sum property). For example, they can be used to assess how a voting reform reshapes the distribution of votes across parties or how an educational intervention changes the allocation of students across majors. However, they have some important limitations.
The first is that treatment often affects not only the relative composition of outcomes but also their total level, the extensive margin. In such cases, focusing solely on shares provides an incomplete picture. A voting reform may reshape the distribution of votes while also changing total turnout. A minimum-wage increase may alter the probabilities of being in different employment categories and affect the overall number of workers. An education reform may shift students across majors while simultaneously changing the total number of graduates. Likewise, a carbon policy may influence the shares of energy sources in electricity production and the total electricity generated. Those methods cannot jointly and consistently recover treatment effects on both shares and totals. Yet, many empirical questions require a framework that coherently incorporates absolute and relative changes simultaneously. The second problem is that those approaches lack economic grounding, as they cannot be supported by behavioral or structural relationships that typically govern how categorical margins evolve. This makes it harder to intuitively evaluate their meaning in practice.
This paper introduces Compositional Difference-in-Differences (CoDiD), a framework for analyzing vectors of category-specific quantities (compositional vectors) in a DiD setting while jointly capturing effects on both the intensive and extensive margins. The key identifying condition is a parallel growth assumption, which delivers point identification of the entire counterfactual compositional vector, and therefore of both counterfactual shares and totals. Parallel growth is simply a parallel trends assumption applied to the log-quantities of each category: in the absence of treatment, every category’s absolute size would expand or contract at the same proportional rate in the treated and control groups. To justify this assumption, I model the observed category quantities as aggregates of individuals’ discrete choices within a standard random utility framework. In this setting, parallel growth is equivalent to assuming parallel trends in the latent expected utilities associated with each category. Intuitively, absent treatment, both groups would experience the same evolution in how attractive each option is on average. For instance, if the expected utility of voting for the Democratic candidate rises in the control group, it would have risen by the same amount in the treated group had treatment not occurred. This interpretation clarifies the causal parameter: the treatment effect captures the extent to which the policy shifts the underlying desirability of each category.
The implied evolution of the shares corresponds to parallel trends in the log-odds transformation of category shares. In the same random utility model, log-odds reflect differences in latent utilities and therefore capture relative preferences. Under this interpretation, parallel growth means that, absent treatment, the evolution of relative preferences between any pair of categories would be the same in both groups. Put differently, the way individuals substitute between categories over time in the control group would mirror the untreated evolution in the treated group. Finally, from a geometric perspective, each share vector corresponds to a point in the simplex, and its evolution traces a path within this constrained space. The parallel growth assumption implies that, absent treatment, the treated group’s counterfactual path would be parallel to the control group’s path under Aitchison geometry, the standard geometric framework for compositional data (Aitchison (1982); Egozcue et al. (2003)).
In settings with multiple time periods, the parallel growth assumption can be evaluated visually by plotting the log-transformed quantities over time. If the pre-treatment trends do not appear approximately parallel, the credibility of the assumption weakens. To address such cases, I propose a relaxation in the spirit of Ban and Kédagni (2022). The idea is to bound the counterfactual log differences between groups by restricting their changes to lie within the convex hull of observed pre-treatment changes. This procedure yields bounds, rather than point estimates, for both the counterfactual categorical quantities and the resulting counterfactual distributions.
As an application, I examine the impact of the early voting program on the 2008 U.S. presidential elections. These programs, allowing voters to cast ballots before election day, either in person or by mail, have been the subject of intense partisan debate. I revisit this question by analyzing both total turnout and the compositional shift in votes for parties. I consider Maryland and New Jersey, which introduced early voting before 2008 in the treated group, while Pennsylvania and New York serve as the control group. As expected, the treatment increased overall turnout ( increase of $4.4\%$) of total number of voters. The ATT shows that the Republican party lost 1.11 percentage points, 0.9 in favor of democrats, and 0.3 in favor of alternative parties.
@@
First, it contributes to the extensive DiD literature. Classical DiD focuses on scalar outcomes, estimating treatment effects on the mean, with surveys and applications summarized in Lechner (2011); Cafri et al. (2019); De Chaisemartin and d’Haultfoeuille (2023); Roth et al. (2023); Baker et al. (2025). Recent works have extended DiD to more complex settings: Callaway and Sant’Anna (2020) develops methods for staggered treatment adoption, Arkhangelsky et al. (2021) combines DiD with synthetic controls, and Manski and Pepper (2018); Rambachan and Roth (2020); Ban and Kédagni (2022) relax the parallel trends assumption. A second wave of research extends DiD to entire outcome distributions. Early contributions include quantile DiD Meyer et al. (1990) and the changes-in-changes (CiC) approach Athey and Imbens (2006), which identifies counterfactual distributions non-parametrically. Subsequent work has generalized these ideas to multivariate outcomes Torous et al. (2021), settings where identifying assumptions apply to cumulative distribution functions Havnes and Mogstad (2015); Roth and Sant’Anna (2021), copula-based approaches Callaway et al. (2018); Callaway and Li (2019); Ghanem et al. (2023), and methods using characteristic functions Bonhomme and Sauder (2011). However, these methods do not directly address categorical outcome distributions, which are shares. Graves et al. (2022) proposes a DiD for categorical outcomes while relying on a linear parallel trends for proportions, which I explained earlier, is not appropriate in general. Zhou et al. (2025) extends the difference-in-differences framework to non-Euclidean data, including shares, by employing Fréchet means and geodesic transport. While their approach is mathematically elegant, the identifying assumptions are primarily geometric and offer limited economic interpretability. In contrast, my framework delivers both counterfactual quantities and shares that retain economic intuition while incorporating a geometric perspective specifically designed for categorical outcomes. Another related paper is Tchetgen Tchetgen et al. (2024), who propose a general DiD framework applicable to count data in the binomial setting. My analysis differs by focusing on the multinomial case, allowing for richer categorical structures and more general forms of distributional change. Second, the paper contributes to the compositional data analysis (CoDA) literature. CoDA, pioneered by Aitchison (1982, 1990, 1992, 2002), provides tools for analyzing data constrained to the simplex. Subsequent developments Egozcue et al. (2003); Billheimer et al. (2001); Barceló-Vidal et al. (2001) refine transformations and models that respect the geometric structure of compositional data. More recently, Arnold et al. (2020) connected CoDA to causal inference. I also combine compositional data methods with econometric identification strategies to analyze causal effects when outcomes are shares, providing a bridge between DiD and compositional statistics.
@@
The paper is organized as follows. Section 1 introduces the canonical $2\times 2$ case, presenting the main identifying assumption along with its economic and geometric interpretations. Section 2 generalizes the framework to settings with multiple time periods. Section 3 illustrates the methodology through the empirical application. Finally, Section 4 concludes.
@@ 1. Canonical $2\times 2$ DiD Case
In this section, I focus on the canonical, two-group, two-time period case in the DiD framework, and I clearly expose the assumptions and their implications. I will start with the analytical framework, later discuss the main identifying assumption, and after all the economic and geometric interpretations.
1.1. Analytical framework
In this section, I introduce the basic framework of my method. Consider some aggregated unit for which we observe the following vector of categorical quantities:
$q=(q_{1},q_{2},\dots,q_{p}).$
@@ 1.1. Analytical framework
I consider a setting with two groups, $g\in\{0,1\}$, where $g=1$ is the treated group and $g=0$ is the control group. These groups are observed in two time periods, $t\in\{0,1\}$, with $t=0$ being the pre-treatment period and $t=1$ the post-treatment period. For any category $k$, I define $q^{0}_{k,g,t}$ as the potential untreated quantity (the quantity in $k$ if the group $g$ were untreated at time $t$). I also have $q^{1}_{k,g,t}$: the potential treated quantity (the quantity in $k$ if the group $g$ were treated at time $t$). For the untreated potential outcomes, I aggregate across categories to define the total quantity for a group-period:
$S^{0}_{g,t}=\sum_{k=1}^{p}q^{0}_{k,g,t}.$
The share (or probability) for category $k$ is then:
$\pi^{0}_{k,g,t}=\frac{q^{0}_{k,g,t}}{S^{0}_{g,t}}.$
The vector of quantities and the entire shares are represented as:
$q^{0}_{g,t}=\left[q^{0}_{1,g,t},\dots,q^{0}_{p,g,t}\right],\quad\pi^{0}_{g,t}=\left[\pi^{0}_{1,g,t},\dots,\pi^{0}_{p,g,t}\right],\quad\text{where}\quad\sum_{k=1}^{p}\pi^{0}_{k,g,t}=1.$
All the same definitions apply to the treated potential quantities, $q^{1}_{k,g,t}$, and their associated totals and shares.
@@
From the data, I can identify the vectors $q^{0}_{0,0}$ for the control group and $q^{0}_{1,0}$ for the treated group in the pre-treatment period. In the post-treatment period, I can identify from data the untreated vector of quantities for the control group, $q^{0}_{0,1}$, and the treated counterpart for the treated group, $q^{1}_{1,1}$. The counterfactual of interest is the vector of quantities that I would have observed for the treated group in the absence of treatment $q^{0}_{1,1}$. Identifying this missing object is the core challenge of my analysis. Once this is identified, I can consider the treatment effect parameters.
Treatment effect parameters:
@@
I introduce treatment effect parameters to capture the impact of the intervention. The first is the growth treatment effect on the treated (GTT), which quantifies the causal effect on the absolute size of a category. For each category $k$ (where $k=1,\ldots,q$), the GTT is defined as the proportional change in its quantity:
$\text{GTT}_{k}=\frac{q^{1}_{k,1,1}}{q^{0}_{k,1,1}}-1,\quad GTT=\frac{S^{1}_{1,1}}{S^{0}_{1,1}}-1.$ (1)
@@
This parameter has an intuitive interpretation: $GTT=0$ implies the treatment had no effect on the absolute. $GTT>0$ indicates that the treatment caused category $k$ to grow in absolute terms. For example, a GTT of $0.15$ means the category’s size increased by $15\%$ due to the treatment. $GTT<0$ signifies that the treatment caused the category to shrink. A GTT of 0.10-0.10, for instance, corresponds to a $10\%$ decline.
When comparing the shares, there are several ways to do it, depending on the type of effects that we would like to emphasize. To quantify the absolute change in shares, we can focus on the Average treatment effect on the treated (ATT), defined as the simple difference in category proportions (or shares) between the treated potential outcome distribution and its untreated benchmark.
$\text{ATT}=\pi_{1,1}^{1}-\pi_{1,1}^{0}.$ (2)
This representation is intuitive, easy to interpret, and directly reflects the absolute reallocation of probability mass across categories induced by the treatment.
@@
Sometimes, it might be interesting to see which category has gained relatively the most as a result of treatment. The compositional treatment effect on the treated (CTT) captures how treatment redistributes weight across categories proportionally and also tells us which category gained relatively the most. To define it, I consider the compositional difference operator commonly used in compositional data analysis to compare shares. This operator quantifies how each category expands or contracts relative to all others. For intuition, fix a category $k$ and consider its proportional change:
@@
$r_{k}=\frac{\pi_{1,1}^{1}(c_{k})}{\pi_{1,1}^{0}(c_{k})},\quad\begin{cases}r_{k}>1&\text{category }k\text{ grew},\\ r_{k}<1&\text{category }k\text{ shrunk},\\ r_{k}=1&\text{no change}.\end{cases}$
@@
Dividing by the total change gives a vector showing each category’s relative gain or loss. I obtain the compositional difference operation $(\ominus)$ used in compositional data analysis for the control group.
$\pi_{1,1}^{1}\ominus\pi_{1,1}^{0}=\frac{1}{\sum_{k=1}^{p}r_{k}}\,\big[r_{1},r_{2},\dots,r_{p}\big]=\frac{1}{\sum_{k=1}^{p}\frac{\pi_{k,1,1}^{1}}{\pi_{k,0,0}^{0}}}\left[\frac{\pi_{1,1,1}^{1}}{\pi_{1,1,1}^{0}},\dots,\frac{\pi_{p,1,1}^{1}}{\pi_{p,1,1}^{0}}\right].$ (3)